You and your sister are saving money from your allowances. you have $25 and sace $3 each week. Your sister has $40 and saves $2 each week. After how many weeks will you and your sister have the same amount of money?

Answer by checkley79(3341)

25+3x=40+2x
3x-2x=40-25
x=15 weeks of savings will result in the same amount of savings.
Proof:
25+3*15=40+2*15
25+45=40+30
70=70